OpenStudy (anonymous):

Heinz has a list of possible functions. Pick one of the g(x) functions below, show how to find the zeros, and then describe to Heinz the other key features of g(x). g(x) = x3 – x2 – 4x + 4 g(x) = x3 + 2x2 – 9x – 18 g(x) = x3 – 3x2 – 4x + 12 g(x) = x3 + 2x2 – 25x – 50 g(x) = 2x3 + 14x2 – 2x – 14

OpenStudy (anonymous):

@nincompoop

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

someone helppppppppppppppppp

OpenStudy (anonymous):

@Awolflover1

OpenStudy (anonymous):

@chrissyC.

OpenStudy (anonymous):

g(x) = x^3 – x^2 – 4x + 4 The zeros are found by factoring the equation. x^2(x-1) - 4(x -1) = 0 (x^2 -4)(x-1) = 0 (x+2)(x-2)(x-1) = 0 x = 2, -2, and 1 The other key features are: g'(x) gives the slope g''(x) gives the concavity g'(x) = 0 will give the critical points g''(x) = 0 will give the point of inflection So A.

OpenStudy (anonymous):

no you pick one then show how to find the zeros

OpenStudy (anonymous):

exactly what i did ^^

OpenStudy (anonymous):

i picked A. and shows which ones show the 0's

OpenStudy (anonymous):

lol ok thx peanut <3

OpenStudy (anonymous):

lml yw jelly Cx <3